C6H6 (l) + Br2 (l) + C6H5Br (l) +HBr (g)


If 54.2 g of benzene is mixed with 122 g of bromine,


Which one is the limiting reagent?

1- Br2

2- C6H6


How many grams of bromobenzene is formed in the reaction?

Mass=..........g




1
Expert's answer
2021-03-01T02:35:00-0500

To get limiting reagent first calculate moles of each reactant.

Mole= mass/molar mass

(Mole)br2= 54.2/160= 0.33

(Mole)benzene= 122/78= 1.56

Hence, mole of Br2 is less so it will be limiting reagent.


Mole of bromobenze= 0.33

Mass of bromobenze= mole×molar mass


Mass of bromobenze= 0.33×157

Mass= 51.8gm